# Design Optimizer for Soft Growing Robot Manipulators in Three-Dimensional Environments

### v1

@@ 3.3 Fitness Evaluation from the Objective Functions
    $f_{2}(\psi,t,h,o)$: minimize the difference between the orientation of the robot’s last link and the desired reaching orientation for each target in the environment – solve the inverse kinematics; and
  3. 3.
    $f_{3}(\psi,t,h,o)$: minimize the robot’s components (i.e., its length, number of links), to promote quicker and more cost-effective production.
(b) Trade-off nodes/orientation
Figure 5: (a) Instead of directly aiming for the only target in the environment, this robot moves in random directions before completing the task, leading to a “wavy” configuration. (b) Trade-off problem between the number of nodes and reaching orientation. On the left, the robot reaches the target with the minimum possible number of nodes (it is assumed that all internal links are at their maximum length within the specified bounds), but it cannot reach the desired orientation. On the right, an additional node ensures that the robot achieves the correct orientation to reach the target.
In multi-objective optimization problems (MOOPs), some or all of the objectives may conflict with each other, meaning that optimizing one objective can adversely affect another. As a result, instead of finding a single solution, we typically obtain a set of trade-off solutions that represent the Pareto optimal front. Solutions on this front are non-dominated by any other solution, meaning no other solution is better across all objectives. However, as trade-off solutions, they may be optimal for one objective but not necessarily for others [77, 78]. In the context of this work, treating all objectives as equally important could result in trade-off solutions that are not desirable for our needs. For instance, while the minimum number of links might be a trivial single-link robot, it could be impractical if the robot cannot reach the target – i.e., a single-link robot cannot steer! Conversely, a robot with an excessive number of useless links (e.g., a “wavy” robot as shown in Fig. 5(a)) is clearly not the optimal solution we seek. Clearly, we are not interested in a robot that fails to reach the target, but we do need to balance the properties defining the robot (e.g., number of links) with its ability to achieve the desired orientation. Minimizing the number of links might compromise the ability to achieve the desired orientation, whereas an additional link can help meet the orientation requirement – as shown in Fig. 5(b). When dealing with multiple targets, achieving the exact desired orientation may be challenging, making trade-offs necessary. However, when feasible, it is preferable to use more links to better satisfy the orientation requirement – thus, $f_{2}$ takes precedence over $f_{3}$, and $f_{1}$ is given overall priority. Based on this, the optimization problem becomes to reach the target’s pose using the least amount of resources. Therefore, we should first verify that the configuration being examined can reach the target ($f_{1}$), and then proceed to minimize its components. It is important to note that although the robot must precisely reach the target, $f_{1}$ should not be treated as a constraint (i.e., solutions with a non-zero tip-to-target distance should not be deemed infeasible). If $f_{1}$ is treated as a constraint, the search algorithm may lose its heuristic effectiveness, leading to random convergence [^4] This claim is debatable, and it depends on the constraint-handling method employed: it would hold true for death or static penalty methods, but it could be valuable to explore results using dynamic methods or separation-based approaches [84], despite the higher risk of obtaining infeasible solutions only..
Using methods that retrieve Pareto optimal solutions, known as generating methods [85], may not be ideal for our situation unless we intentionally limit part of the objective space. Instead, we aim to integrate the three objectives into a single function, referred to as a fitness function. In addition to prioritizing the first objective, single-objective optimization methods generally have lower computational complexities than multi-objective methods.
To solve this design optimization problem, we employ the Rank Partitioning method, which was introduced in the optimizer for the 2D case [67]. Rank partitioning is a non-parametric survival strategy for Multi-Objective EC algorithms (or generically for population-based optimization methods) that ranks each individual based on a predefined order of the objectives. It sorts the population for the first objective (i.e., the one with the highest priority) and then creates partitions based on their objective rank; then, it sorts each partition for the second objective, and so on. Finally, it returns only one solution: the optimal based on the given order of priorities. It can be applied to multi or many-objective optimization (i.e., with more than three objectives [86]).
3.3.1 Solving the Inverse Kinematics with the Desired Orientation
The first two objectives are to ensure that the end effector of each configuration reaches the respective target ($f_{1}$) with the desired orientation ($f_{2}$). This is a well-known problem of inverse kinematics [87], which can be solved by minimizing the distance between the end effector and the target for each configuration. This is usually solved in task space (i.e., Cartesian coordinates in the environment) rather than in joint space (i.e., polar coordinates defining the values of each robotic joint). Therefore, since the configurations $\zeta$ in Eqn. (1) are defined in polar coordinates ($\zeta^{p}$), it is necessary to transform it into its corresponding configuration in Cartesian coordinates ($\zeta^{c}$) through forward kinematics as in Eqn. (5) – details are reported in Section 1.1 of the supplementary material. The configuration $\zeta^{c}$ is composed of $n+1$ points – i.e., all nodes plus the end effector.
@@ 3.3.1 Solving the Inverse Kinematics with the Desired Orientation
Figure 6: The phenotype of a robot generated with the genotype described in Eqn. (2), having $n=10$ links. In this case, only five are actually active to reach the target correctly ($\bar{n}=5$). Moreover, the last link is not fully extended, unlike the internal links that make up the rest of the body: this information is codified by extending the genotype as in Eqn. (6). The five links after the end effector (represented by dotted lines) do not need to be included in the design $\delta$.
@@
Figure 7: (a) When the inverse kinematic solver minimizes the distance node-target, there is no guarantee for the robot to have the last link aligned with the desired target orientation; however, when the solver minimizes the distance node-orientation, the robot will reach the desired orientation and can simply grow straight to the target after that. (b) The phenotype codified by a solution $\psi$ ignores every link after $\epsilon$ originally generated in the genotype and aims directly to the target for the remaining $\bar{n}-\epsilon$ links. The last link might not be fully everted, with a length smaller than the one originally generated in the genotype – indicated with the term $l_{\bar{n}}$.
The first issue to be addressed is the location of the end effector in a configuration: the population of any EC algorithm will be filled with random values for each solution $\psi$, meaning that the robotic configurations will explore the environment randomly to eventually converge with the evolutionary metaheuristic. Fig. 6 shows that the tip of the robot, or the end effector, can be placed anywhere on any link as the robot grows along its configuration. The link whose endpoint corresponds to the end effector is indicated with the index of $\bar{n}$. In case $n\neq\bar{n}$, in which $\bar{n}$ links are active and the rest $n-\bar{n}$ are folded within the robot, it is possible that the $\bar{n}$-th link is only partially everted – i.e., the target is reached before the link is fully grown. Based on the definition in Eqn. (2), multiple configurations share the same design because the robot can take different configurations with different angles, but the link lengths cannot change. However, when different targets require a different amount of growth, as shown in Fig. 4(b), it is necessary to modify the value of the last link for each configuration (i.e., $\bar{n}$) such that the link length only takes into account the active part of the link. In that case, the $\bar{n}$-th link length is not taken from the design $\delta$ (i.e., the last row of $\psi$ in Eqn. (2)), but is defined as the distance between the $\bar{n}$-th node and the target, and indicated with the term $l_{\bar{n}}$.
The second issue to be addressed is to guarantee that the configuration’s last link(s) lie(s) on the target’s desired orientation segment. The left-hand side of Fig. 7(a) shows that when minimizing the distance node $\bar{n}$ to target, the optimized configuration might end up not being aligned with the desired orientation. Rather than also minimizing the difference between the orientation of the robot’s last link and the desired one, it is more efficient to combine both objectives $f_{1}$ and $f_{2}$ into a single one, namely $f_{1-2}$. The right-hand side of Fig. 7(a) shows an alternative minimizing metric: reducing the distance between the end effector $\bar{n}+1$ to the target’s desired orientation segment. With this formulation, once the distance is minimized, the robotic configuration will intersect with the target’s orientation segment at the end effector $\bar{n}+1$. However, the end effector would still not be on the target, but somewhere on the segment: at this point, the robot can simply evert toward the target to eventually reach it – as shown in Fig. 7(b). This requires additional information to be stored in the genotype of a solution $\psi$. Firstly, the node selected to reach the target’s orientation segment is not the end effector but the closest node to the segment (at index $\epsilon$) among all the configuration nodes. Secondly, the rotation angles at the $\epsilon$-th joint (namely, $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$) must not follow the ones indicated in the configuration $\zeta$, but should be geometrically calculated such that the robot is oriented in the direction of the target (the procedure is reported in Sections 1.2 and 1.3 of the supplementary material). Fig. 7(b) shows the phenotype described by the genotype with the new extra information.
Considering these two issues, the additional information needs to be codified in the data structure/genotype of solutions $\psi$, whose elements define the decision variables of the optimization problem. The matrix $\psi$ then is defined as in Eqn. (6), where the left submatrix is the solution genotype $\psi$ defined in Eqn. (2), and the right submatrix is composed of the extra information for each configuration. Specifically, for each configuration, every pair of rows of the right submatrix contains:
@@
    $\epsilon$, the index of the closest node to the target’s orientation segment;
@@
    $\theta^{(x)}_{\epsilon}$, the angle on the x-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
@@
    $\theta^{(y)}_{\epsilon}$, the angle on the y-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
@@
    $\bar{n}$, the number of active nodes/links for each configuration (with $\bar{n}\leq n$) – namely the index of the last active node of a configuration; and
  • 5.
    $l_{\bar{n}}$, the length of the last link for each configuration, right before the end effector.
Note that, to keep a matrix shape, some elements are set to zero as they are unused: (i) the second row of each configuration, which defines rotations on the y-axis, only contains the last angle rotation $\theta^{(y)}_{\epsilon}$; and (ii) the last row, which defines the link lengths, are all fixed for a single design and do not need any extra information.
@@
While the left submatrix is either randomly generated in the first generation of the EC algorithm or generated by variation when producing offspring, the right submatrix is filled during solution evaluation based on the geometrical requirements needed to reach each specific target. Also, it is important to note that the link lengths obtained in the design $\delta$ in Eqn. (3) serve as a reference only for the target farthest from the home base $h$, and that different configurations may use partial information from $\delta$ based on the number of active links.
The combined fitness $f_{1-2}$ is expressed as in Eqn. (7), where the distance between a node of the robot and the segment $s_{i}$ is calculated with the point-segment distance (introduced by Stroppa et al. [88] for 2D and extended for this work’s purposes in 3D as reported in Algorithm S.3 of the supplementary material). This computes the distance from the point to its projection onto the segment or to the nearest edge. This formulation also retrieves the index $\epsilon$ of the node that will intersect with the target’s orientation segment.
3.3.2 Optimizing the Kinematic Components of the Robot
This part is the same as the 2D case presented in the previous work [67], except for having two rotational angles for each node; however, we report it fully for a comprehensive description and reproducibility. The objective functions $f_{1}$ and $f_{2}$ (inverse kinematics) take precedence over $f_{3}$ (optimal components). After combining the first two into a single fitness function $f_{1-2}$, the next step is penalizing excessively long solutions with too many nodes. Ideally, we aim for a robot that gradually aligns with the target’s orientation segment, creating a smooth curve. The robot depicted in the right-hand side of Fig. 5(b) is a better solution compared to the one shown in Fig. 5(a): it not only positions all nodes closer to the orientation segment but also uses fewer nodes to reach the target.
Objective $f_{3}$ focuses on optimizing the robot’s components to minimize manufacturing time and resources and cannot be captured by a single property. Thus, we treat it as a macro objective composed of the following:
  • 1.
    $f_{3.1}$: minimize the number of links of each configuration to reduce redundant DoFs, as in Eqn. (8);
    $f_{3.1}=\sum_{i=1}^{|t|}\bar{n}_{i}$ (8)
  • 2.
    $f_{3.2}$: minimize the overall length of the robot to promote shorter configurations, as in Eqn. (9); and
  • 3.
    $f_{3.3}$: the change in direction between links in all configurations, or undulation, to favor trajectories that are non-intersecting and not “wavy”, calculated as a percentage of the total length of all configurations, as shown in Eqn. (10) – note that the formula checks the change in the angle between the $j$-th and $(j+1)$-th link in both the $xz$ and $yz$-plane projections ($\Delta\theta^{(xz)}$ and $\Delta\theta^{(yz)}$, respectively).
The priorities among these three sub-objectives can be adjusted to refine our preference-based method, potentially eliminating the need for weight parameters. Although it might appear that the number of links and the total length of the robot ($f_{3.1}$ and $f_{3.2}$, respectively) have a linear relationship (i.e., more links lead to a longer robot), this is not necessarily true. A robot with many small links can be shorter than one with fewer, very long links. We prefer robots with fewer links because the time required to manufacture a link is not affected by its length ($f_{3.1}$ takes precedence over $f_{3.2}$). On the other hand, the number of links and their orientations ($f_{3.1}$ and $f_{3.3}$, respectively) affect the robot’s shape differently. A robot with many links is more likely to be wavy, as illustrated in Fig. 5(a). Whether high-DoF non-wavy robots are preferable to low-DoF wavy ones may not be straightforward and could depend on the specific task (e.g., navigating a maze might require varying directions); however, we found that fewer links are preferable for manufacturing ($f_{3.1}$ takes precedence over $f_{3.3}$), but other manufacturers might have different priorities and choose to reverse this order. Finally, the undulation of a robot and its length ($f_{3.3}$ and $f_{3.2}$, respectively) are strongly correlated. A robot that directly targets the destination will generally be shorter than one that moves through the environment before reaching the target, as demonstrated in Fig. 5(a). Therefore, we consider it more important to construct non-wavy robots, which will ultimately lead to shorter configurations ($f_{3.3}$ takes precedence over $f_{3.2}$).
Furthermore, we found that to enhance the quality of the solutions obtained, the objective $f_{3.1}$ should be broken down into two sub-objectives:
  • 1.
    $f_{3.1a}$: minimize the number of links to reach the target’s orientation segment, as in Eqn. (11); and
    $f_{3.1a}=\sum_{i=1}^{|t|}(\epsilon_{i}-1)$ (11)
  • 2.
    $f_{3.1b}$: positioned along the target’s orientation segment, as in Eqn. (12).
    $f_{3.1b}=\sum_{i=1}^{|t|}(\bar{n}_{i}-\left(\epsilon_{i}-1\right))$ (12)
By making this division, we specifically direct how the configuration should be arranged rather than just minimizing the total number of links across different configurations. This involves prioritizing the reduction of links required to align with the target’s orientation segment. Once this segment is reached, we then focus on minimizing the number of links needed to reach the target itself. Note that the final step can be effectively achieved by manually maximizing the lengths of these links. Although the optimization algorithm should ideally learn this heuristic, practical implementation might require some manual adjustments due to the extensive search space.
3.4 Separation of the Objectives with Rank Partitioning
1
input : $N$, population size.
input : $F$, a matrix $N\times 9$ having the following information for each individual: fitness for each objective (integer values first followed by real values), fitness after partitioning ($rank$), reference to the individual in the population ($ref$). A single row of $F$ is: [(1)[(1)[ ( 1 ): $F_{1-2\mathbb{N}}$, $(2)$: $F_{3.1a}$, $(3)$: $F_{3.3}$, $(4)$: $F_{3.1b}$, $(5)$: $F_{3.2\mathbb{N}}$, $(6)$: $F_{1-2\mathbb{R}}$, $(7)$: $F_{3.2\mathbb{R}}$, $(8)$: $rank$, $(9)$: $ref]_{i}$ $\forall i\in[1,N]$.
input : $binSize$, distance between partitions for each real-coded objective (i.e., one for $F_{1-2}$ and one for $F_{3.2}$).
output : The updated matrix $F$ with the value of $rank$ after partitioning.
2
3begin
4       for $i\leftarrow 1$ to $N$ do
5             $F_{i,1}\leftarrow F_{i,6}-\textnormal{{mod}}(F_{i,6},binSize_{1-2})$
6             $F_{i,5}\leftarrow F_{i,7}-\textnormal{{mod}}(F_{i,7},binSize_{3.2})$
7
8
9      sortrows($f,>$) on column 1-5 for all rows
10       $d\leftarrow 0_{N\times 5}$
11       for $i\leftarrow 2$ to $N$ do
12             for $j\leftarrow 1$ to $5$ do
13                   $d_{i,j}\leftarrow\textnormal{{abs}}(F_{i,j}-F_{i-1,j})$
14
15
16
17      $start\leftarrow 1$
18       for $i\leftarrow 1$ to $N$ do
19             for $j\leftarrow 1$ to $5$ do
20                   $partLimit\leftarrow partLimit+d_{i,j}$
21
22            if $partLimit>0$ then
23                   $end\leftarrow i$
24                   sortrows($f,>$) on column 6-7 from row start to end-1
25                   $start\leftarrow end$
26
27
28
29      for $i\leftarrow 1$ to $N$ do
30             $F_{i,8}\leftarrow i$
31
32      return $F$
33
Algorithm 1 Rank Partitioning [67]
Based on the observations about objective priorities, we employ a strategy that prioritizes solutions with fewer, well-structured components, but only if they are sufficiently close to the target’s orientation segment. The Rank Partitioning algorithm was introduced in the previous work by Stroppa [67] and is used to retrieve the best solution of a MOOP given a predetermined order of priorities among objectives, without the burden of compromising the optimality of the highest prioritized objectives and choosing weights for each of them (i.e., unlike weighted sum methods [78, 89]).
Rank partitioning divides the population into $k$ partitions (bins) with fixed intervals (bin sizes) according to their kinematic fitness $f_{1-2}$. Individuals with better fitness are assigned to the first segments, while those with worse fitness are placed in the last segments – in this specific context, where all objectives will be minimized, the best fitness is the one with the lowest value. This is repeated for each objective: the partitions made for the first objective will be divided into more partitions based on the second objective, and so on, until all the objectives are partitioned. The key idea is that individuals within the same partition are considered equally optimal in terms of the partitioned objective so that the best among them can be selected based on the following objectives. Just like in the 2D scenario [67], we applied Rank Partitioning in the following order for each partition:
  1. 1.
    distance from the target/inverse kinematics ($f_{1-2}$);
  2. 2.
    number of links to reach the targets’ orientation segment ($f_{3.1a}$);
  3. 3.
    percentage of undulation among configurations ($f_{3.3}$);
  4. 4.
    number of links on the targets’ orientation segment ($f_{3.1b}$); and
  5. 5.
    length of the robot/final design ($f_{3.2}$).
This approach works only when ties happen within partitions, leading to nested sub-partitions that follow the prioritized order of objectives. This creates a challenge when handling objectives with continuous domains, as we deal with ranges or bins of fixed intervals rather than individual values (similar to a histogram). To address the issue, Rank Partitioning duplicates all continuous objectives (i.e., $f_{1-2}$ and $f_{3.2}$) by adding a further objective composed by the lower-bound discrete value of their bin. Discrete objectives can be used with no modification (i.e., $f_{3.1}\in[1,n\cdot|t|]$ and $f_{3.3}\in[0,100]$). In this manner, the partitioning operation can be resolved using a straightforward sorting algorithm that accounts for ties. When a tie occurs in the objective with the lowest priority, individuals are ranked based on their original continuous values of continuous-domain objectives (i.e., $f_{1-2}$ and $f_{3.2}$), which should be preserved rather than substituted by their discrete bin values. Finally, Rank Partitioning updates the fitness value by assigning a rank to each individual in the population, which increases linearly according to their position within the partitions (i.e., sequentially numbering each individual after the partitioning process from 1 to $N$, with $N$ being the population size).
Algorithm 1 shows Rank Partitioning on the soft growing robot problem, which runs in $\mathcal{O}(N\texttt{log}N)$, where $N$ represents the population size. The input $F$ is a “fitness” matrix $N\times 9$ containing, for each individual, (i) its kinematic fitness $f_{1-2}$, (ii) the four sub-objectives describing the components $f_{3}$, (iii) the original values of the continuous objectives that were evaluated by their lower-bound bin value, (iv) the updated rank value, calculated by the procedure, will be utilized by the selection operators for the reproduction and survival of individuals, and (v) a reference to the individual in the population [^5] The reference is used to link fitness in the matrix and the corresponding individual in the population since only the fitness matrix will be sorted to save computational time. This is simply a pointer to the respective individual in the population.. Lines 2-4 convert continuous objectives (denoted by the subscript $\mathbb{R}$) into discrete values (denoted by subscripts $\mathbb{N}$) based on the specified bin size [^6] Note that the subscript $\mathbb{N}$ is just a placeholder since these values are not limited to natural numbers only as their bin size can be smaller than 1.. Line 5 sorts the rows of the fitness matrix according to the five discrete objectives in the order provided. Lines 6-9 determine partition boundaries by computing differences between consecutive elements in each sorted column of the fitness matrix: if two consecutive elements in the same column are equal, they belong to the same partition. Summing these differences for each row (lines 12-13) helps the algorithm identify partition boundaries. Lines 11-17 re-sort each row by the original continuous values, but only within partitions, ensuring the order from line 5 remains intact. Finally, lines 18-20 assign the updated rank to each individual.
In multi-objective optimization, this approach acts as a restriction within the objective space, focusing only on the segment of the Pareto optimal front with the lowest values for one objective: the first partition represents the optimal solution for $f_{1-2}$, allowing us to select the individual that minimizes $f_{3}$ as the best fit. The key benefits of this method are: (i) it enforces priority on a specific objective (e.g., inverse kinematics in our case), (ii) it eliminates the need to select weight values as in typical preference-based methods, and (iii) it provides flexibility for future research to incorporate additional sub-objectives that may be relevant but were not considered in this study (i.e., for tackling many-objective optimization problems).

### v2

@@ 3.2 Fitness Evaluation and Dynamic Genotype Change
    $f_{2}$: For each configuration, minimize the displacement between the desired orientation segment and the robot’s last link – inverse kinematic solver.
    This is done by transforming configurations $\zeta$ in Eqn. (1) from polar coordinates ($\zeta^{p}$, i.e., in joint space) to Cartesian coordinates ($\zeta^{c}$, i.e., in task space) through forward kinematics as in Eqn. (4) – details are reported in Section 1.1 of the Supplementary Material. The configuration $\zeta^{c}$ includes the end effector, resulting in $n+1$ nodes.
@@
    • (a)
      $f_{1-2}$: We merged these two objectives into a single one, which is to minimize the distance between the closest node to the target’s orientation segment and the segment itself, allowing the robot to correctly aim for the target’s pose as shown in Fig. 5(a). This is calculated according to Eqn. (5) – the distance from a node to the $i$-th target’s orientation segment $s_{i}$ is reported in Algorithm S.3 of the Supplementary Material.
      This formulation requires some additional information to the basic genotype described in Eqn. (2): once the closest node to the target’s orientation segment is found (at index $\epsilon$), the remainder of the configuration’s link is unused (the dotted links in Fig. 5(a)), and we replace them with links that go straight to the target. This results in a different orientation from the angles specified for the $\epsilon$-th joint (defined as $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$), a number of actually used links in the overall configuration ($\bar{n}$), and the length of the last link that does not necessarily need to be fully unfolded ($l_{\bar{n}}$). The procedures for retrieving these values are reported in the Supplementary Material.
@@
    Figure 5: (a) When the inverse kinematic solver minimizes the distance node-target $d$, there is no guarantee for the robot to have the last link aligned with the desired target orientation; however, when the solver minimizes the distance node-orientation, the robot will reach the desired orientation and can simply grow straight to the target after that. (b) The phenotype codified by a solution $\psi$ ignores every link after $\epsilon$ originally generated in the genotype and aims directly to the target for the remaining $\bar{n}-\epsilon$ links. The last link might not be fully everted, with a length smaller than the one originally generated in the genotype – indicated with the term $l_{\bar{n}}$.
  • 3.
    $f_{3}$: For each configuration, minimize the robot’s components to promote a quicker and more cost-effective manufacture. We split this objective into the following:
    • (a)
      $f_{3.1a}$: For each configuration, minimize the number of links needed to reach the target’s orientation segment, as in Eqn. (6).
      $f_{3.1a}=\sum_{i=1}^{|t|}(\epsilon_{i}-1)$ (6)
    • (b)
      $f_{3.1b}$: For each configuration, minimize the number of links lying on the target’s orientation segment, as in Eqn. (7).
      $f_{3.1b}=\sum_{i=1}^{|t|}(\bar{n}_{i}-\left(\epsilon_{i}-1\right))$ (7)
    • (c)
      $f_{3.2}$: Minimize the overall length of the robot, which is the longest robot’s length among all configurations, as in Eqn. (8).
      (a) Wavy robot
      (b) Trade-off nodes/orientation
      Figure 6: (a) A poorly designed robot featuring a wavy configuration. (b) Reducing the number of links and increasing their length helps solve this issue, a trade-off is necessary to make sure the robot smoothly reaches the target.
    • (d)
      $f_{3.3}$: For each configuration, minimize the directional change of the links, or undulation, to avoid “wavy” configurations are shown in Fig. 6. This is evaluated in percentage over the number of links, as shown in Eqn. (9) – note that the formula checks the change in the angle between the $j$-th and $(j+1)$-th link in both the $xz$ and $yz$-plane projections ($\Delta\theta^{(xz)}$ and $\Delta\theta^{(yz)}$, respectively).
Since this evaluation dynamically adds extra information to the genotype, it is necessary to extend the solution matrix $\psi$ by including, for each configuration:
@@
    $\epsilon$, the index of the configuration node/joint that is closest to the target’s orientation segment;
@@
    $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$, the angles on the x and y-axes defining the orientation for the $\epsilon$-th link, dynamically set to go in the direction of the target;
@@
    $\bar{n}$, the number of active nodes/links (with $\bar{n}\leq n$) – namely the index of the last active node of a configuration; and
@@
    $l_{\bar{n}}$, the last link’s length.
The complete matrix $\psi$ is defined as in Eqn. (10), and the respective phenotype that it represents is shown in Fig. 5(b) – example reports one configuration only for the sake of clarity. Note that, being a matrix, some elements are set to zero as they are unused: (i) the second row of each configuration, which defines rotations on the y-axis, only contains the last angle rotation $\theta^{(y)}_{\epsilon}$; and (ii) the last row, which defines the lengths of the links, is fixed for a design and do not need any extra information.
@@
Figure 7: The genotype described in Eqn. (2) yields the phenotype in the example, with a budget of $n=10$ links. However, only five of them are required to correctly reach the target ($\bar{n}=5$), an information included in the genotype of Eqn. (10).